What volume of 0.15 M HCl(aq) must be diluted to make 2.0 L of 0.050 M HCl(aq)?.


1
Expert's answer
2021-02-04T07:07:52-0500

find the formula:

"С\times V=C1\times V1"

"0.15\times V=0.050\times 2"

"V=\frac{0.050\times2}{0.15}"

V=0.666 L